PepsiCo, Inc. (Public, NYSE:PEP). Add to Portfolio. Discuss PEP n Open: 69.56 Mkt Cap: 109.438 P/E: 19.60 09.30 High: 70.20 5Zoomid 5d im 3m 6m YTR 1X Sy 10 Mar Feb 20, 2009 - May 19, 2009 -3.11 (-4,37%) Mar 2008 Apr 2008 May 2008

The above screen shot from Google Finance shows basic stock information for PepsiCo. If you owned 2600 shares of PepsiCo for the period​ shown, how much would you have earned in dividend​ payments?

A- 780

B-1,069.96

C-950.30

D- 1,118
Answer #1

Option D:1118

Given

Number of share N=2600

dividend Payment D=0.43

So Total Dividend =N*D=2600*0.43=1118